A boat is moored to a dock using two ropes exerting force of 400N each on the boat. There is a wind blowing off-shore and keeping the boat away from the dock. What force does the boat experience because of the wind? The two ropes and deck form an isoceles triangle with the angle of 28º at the apex.
1
Expert's answer
2020-05-08T16:11:57-0400
"F=2T\sin{(0.5\cdot28)}"

"F=2(400)\sin{14}=194\ N"